# Robust Self-calibration of Focal Lengths from the Fundamental Matrix

### v2

@@ 1 Introduction
Camera calibration, i.e., the estimation of camera intrinsic parameters, is a fundamental problem in computer vision with many applications. The precision of estimated intrinsic parameters, such as focal length and principal point, significantly affects the precision of tasks in structure-from-motion (SfM) [38],visual localization [33], 3D object detection [45], augmented reality [6], and other applications [39].
Classical calibration methods [46] use known calibration patterns, e.g. checkerboards, or additional knowledge of the observed scene to estimate the camera intrinsics. As such, they are often impractical. On the other hand, self-calibration methods do not require any additional knowledge of the scene geometry and rely on automatic detection of image features in input cameras. Thus, they are preferable in many applications.
@@
The geometry of two uncalibrated cameras is captured by the fundamental matrix [28]. In 1998 Bougnoux [2] showed that assuming square pixels and known principal points of both cameras, it is possible to obtain the focal lengths of the two cameras from the fundamental matrix using a closed-form formula. This provides a convenient way to determine the focal length, since the principal point can usually be assumed to lie in the image center. However, the precision of the estimated focal lengths using the Bougnoux formula [2] and other similar methods [18, 29, 10] is often marred by inaccuracies, sometimes even resulting in physically implausible imaginary focal lengths. This stems from singularities and susceptibility of these methods to noise in fundamental matrices and assumed positions of principal points. Moreover, the singularities occur in very common camera configurations when the principal axes of the cameras are coplanar (i.e. they intersect), which is common when images of a single object of interest are taken from different views.
An alternative approach [14] relies on iterative optimization of a multiterm loss, allowing one to estimate both the focal lengths of the cameras as well as their principal points. This approach addresses some of the drawbacks of the aforementioned methods but still relies on the Bougnoux formula [2] within the iterative process. It is thus susceptible to inaccuracies, especially in the aforementioned degenerate camera configurations.
Similarly to [14], we formulate the problem of estimating focal lengths and principal points from a given fundamental matrix as a constrained optimization problem using priors. Constraints ensure that the solution satisfies the Kruppa equations [27]. To solve the constrained optimization problem, we search for all stationary points of its Lagrangian. This results in a complex system of polynomial equations that we efficiently solve in an iterative way. Each iteration in our formulation involves solving a system of two polynomial equations of degree four in two variables, which is efficiently solved using the Gröbner basis method [23].
Existing approaches, as well as our method, are based on fundamental matrices that can be estimated using different variants of RANSAC [11, 1, 8, 9, 16, 7]. Some of the matrices produced by the 7-point algorithm [15] within RANSAC lead to imaginary focal lengths when decomposed using the Bougnoux formula [2]. We observed that such matrices seldom lead to the final model. Based on this insight, we propose to perform a degeneracy check for such matrices and reject them, thus saving computational time on their further processing, such as scoring and local optimization.
@@
  • A novel iterative method for focal length and principal point estimation from fundamental matrices. Evaluation on synthetic and large-scale real-world datasets [17, 47] shows that our method results in superior accuracy of estimated focal lengths as well as camera poses.
  • A simple and computationally efficient degeneracy check for fundamental matrices leading to imaginary focal lengths within RANSAC. Performance on real-world datasets [17, 47] shows that performing this check and rejecting such models within various RANSAC variants [1, 8, 9, 7] and implementations [24, 3, 16] leads to faster computation and more accurate pose and focal length estimates.
@@ 2.1 Direct Methods
Bougnoux [2] showed that under the assumption of square pixels and known positions of the principal points, it is possible to calculate the focal lengths of both cameras from the fundamental matrix in closed form. Kanatani and Matsunuga [18] derived a closed form solution directly from the elements of the fundamental matrix, avoiding the intermediate epipole computation required by the Bougnoux formula. Melanitis and Maragos [29] formulated a linear system, the solution of which provides the focal lengths.
@@
Similar formulas have been derived for the case when the focal lengths of cameras are known to be equal [41, 18, 4]. In case of equal focal lengths, it is possible to avoid estimating the full 7-DoF fundamental matrix and instead employ a 6-point minimal solver [40] within RANSAC to find the relative poses of cameras along with the focal length. These methods also suffer from generic singularities when the principal axes are parallel, or they intersect, and the camera centers are equidistant from the intersection point [42].
An interesting case is when only one of the focal lengths is unknown. The focal length can be estimated in a closed form solution from the fundamental matrix [44] or using a 6-point minimal solver [5]. A degenerate case for these methods only occurs when the center of the calibrated camera lies on the principal axis of the uncalibrated camera, a situation rarely occurring in practice.
Since focal lengths appear in squared form in the Kruppa equations [27], all the previously mentioned direct methods provide only the squares of focal lengths. This may result in physically implausible imaginary focal lengths due to noise. Depending on the level of noise and camera configuration, imaginary focal lengths may occur at a significant rate (over 20% of samples) [19]. Imaginary focal lengths are less likely to occur when self-calibration is performed using three views [20].
@@ 2.2 Iterative Methods
Hartley and Silpa-Anan [14] pointed out that errors in the assumed positions of principal points may significantly affect the focal lengths estimated by the Bougnoux formula. To remedy this issue, they proposed an iterative method based on Levenberg-Marquardt optimization [25]. During the optimization the fundamental matrix, along with the positions of principal points in both images, are optimized to minimize a multiterm loss. The loss consists of the Sampson error on the correspondences and the distance of the principal points and squared focal lengths from predetermined priors. To prevent imaginary focal lengths, the loss also contains a penalization term which dramatically increases the loss whenever the square focal lengths are below a given threshold. During the optimization procedure, the focal lengths are estimated from the fundamental matrix using the Bougnoux formula. This approach overcomes many of the issues of using the formula alone but may still suffer in the vicinity of degenerate camera configurations, as it still relies on the formula to estimate the focal lengths. Additionally, the procedure is computationally expensive as it requires computation of the Sampson error for all correspondences in each iteration.
Iterative methods can also be used to estimate the camera intrinsics using two or more views [12, 31, 27, 10]. When considering only two views with two unknown focal lengths, these methods should converge to the Bougnoux formula. However, these methods may result in different outcomes in the vicinity of degenerate configurations. For example, the method by Fetzer et al. [10] optimizes for an energy functional based on the Kruppa equations [27]. In degenerate configurations, the functional does not have a single minimum, yet the optimization procedure will be terminated at some point, providing an output different from the one provided by the Bougnoux formula.
@@ 3 Robust focal length estimation
In this paper, similar to Hartley [14], we formulate the problem of estimating focal lengths from a given fundamental matrix $\mathbf{F}$ as an optimization problem using priors.
@@
For fixed principal points, e.g., $\mathbf{c}_{i}=[0,0]$, $i=1,2$, there is a closed form solution for the focal lengths extracted from the given fundamenatal matrix $\mathbf{F}$, i.e., the Bougnoux formula [2]. Since this formula can result in unstable or imaginary estimates, in our formulation, we allow moving the estimated principal points to avoid such instabilities. For non-fixed principal points $\mathbf{c}_{i}$, there are infinitely many decompositions of $\mathbf{F}$ into the essential matrix $\mathbf{E}$ and two calibration matrices of the form (1). Thus, we formulate the problem of estimating focal lengths from $\mathbf{F}$ as a constrained optimization, where during the optimization, the focal lengths, along with the positions of principal points in both images are optimized to minimize the following cost function:
@@
[…]
@@
where $f_{i}^{p}$ and $\mathbf{c}_{i}^{p}=[u_{i}^{p},v_{i}^{p}]^{\top}$, $i=1,2$ are the priors for the focal lengths and the principal points, $w_{i}^{f}$ and $w_{i}^{\mathbf{c}}$ are predetermined weights, and $\kappa_{1}=0$ and $\kappa_{2}=0$ are two Kruppa equations [27] [^2] Note, that there are three Kruppa equations, but only two are linearly independent., which are functions of both focal lengths and principal points that appear in $\omega^{*}_{i}=\mathbf{K}_{i}\mathbf{K}_{i}^{\top}$. The Kruppa equations ensure that for the input fundamental matrix $\mathbf{F}$ and the estimated calibration matrices $\mathbf{K}_{1}$ and $\mathbf{K}_{2}$ (1), the matrix $\mathbf{K}_{1}^{\top}\mathbf{F}\mathbf{K}_{2}$ is a valid essential matrix.
The cost function in (2) has several advantages over the cost function used in [14]. (1) It does not contain a term with the Sampson error on correspondences, which significantly slows down the optimization in [14]. (2) It operates directly on focal lengths $f_{i}$ instead of squared focal lengths $f_{i}^{2}$. (3) The minimization of (2) does not require computation or initialization using the Bougnoux formula [2] as used in [14]. Such an initialization can be very far from the ground truth focal lengths and in many scenarios can result in $f_{i}^{2}\leq 0$. Therefore, minimization of (2) does not need a special penalty term for $f_{i}^{2}\leq 0$ as used in [14].
@@
where $\lambda_{1}$ and $\lambda_{2}$ are the Lagrange multipliers. The constant ‘$-2$’ is introduced for easier subsequent manipulation of the equations and it does not influence the final solution.
@@
Unfortunately, the system of eight equations (4)-(7) is too complex to be efficiently solved using an algebraic method, e.g. using the Gröbner basis method [21, 23]. Next, we propose an iterative method to efficiently solve this system of polynomial equations. Note that in our method, we estimate all stationary points of the Lagrange function $L$ (3). Thus, we avoid potential problems of numerical methods that directly minimize (3). Such methods usually require good initialization and can get stuck in local minima.
@@ 3.1 Iterative method
The proposed iterative solution to equations (4)–(7) follows the idea of iterative triangulation methods [26, 22]. First, let $s^{k-1}=\langle f_{1}^{k-1},f_{2}^{k-1},\mathbf{c}_{1}^{k-1},\mathbf{c}_{2}^{k-1}\rangle$ denote the best current estimate of focal lengths and principal points after the $(k-1)^{th}$ iteration. The prior values $f_{i}^{0}\equiv f_{i}^{p}$, $\mathbf{c}_{i}^{0}\equiv\mathbf{c}_{i}^{p}$, $i=1,2$ are used as an initialization. In the $k^{th}$ iteration the updated estimates $s^{k}=\langle f_{1}^{k},f_{2}^{k},\mathbf{c}_{1}^{k},\mathbf{c}_{2}^{k}\rangle$ are obtained by replacing the focal lengths and principal points $\langle f_{1},f_{2},\mathbf{c}_{1},\mathbf{c}_{2}\rangle$ on the right-hand side of equations (8)–(9) by the current best estimates $s^{k-1}=\langle f_{1}^{k-1},f_{2}^{k-1},\mathbf{c}_{1}^{k-1},\mathbf{c}_{2}^{k-1}\rangle$. In this way, we obtain the expressions for $\Delta f_{i}=\frac{1}{w_{i}^{f}}(\lambda_{1}\frac{\partial\kappa_{1}}{\partial f_{i}}(s^{k-1})+\lambda_{2}\frac{\partial\kappa_{2}}{\partial f_{i}}(s^{k-1}))$ and $\Delta\mathbf{c}_{i}=\frac{1}{w_{i}^{\mathbf{c}}}(\lambda_{1}(\frac{\partial\kappa_{1}}{\partial\mathbf{c}_{i}}(s^{k-1}))^{\top}+\lambda_{2}(\frac{\partial\kappa_{2}}{\partial\mathbf{c}_{i}}(s^{k-1}))^{\top})$, $i=1,2$. Here $\frac{\partial\kappa_{j}}{\partial f_{i}}(s^{k-1})$ and $\frac{\partial\kappa_{j}}{\partial\mathbf{c}_{i}}(s^{k-1})$, $i,j=1,2$ are partial derivatives of $\kappa_{j}$ evaluated at $s^{k-1}=\langle f_{1}^{k-1},f_{2}^{k-1},\mathbf{c}_{1}^{k-1},\mathbf{c}_{2}^{k-1}\rangle$. After this substitution, $\Delta f_{i}$ and $\Delta\mathbf{c}_{i}$ are functions of two Lagrange multipliers $\lambda_{1}$ and $\lambda_{2}$.
@@
The updated Kruppa equations (6) and (7), which we denote $\kappa_{1}^{k}=0$ and $\kappa_{2}^{k}=0$, form a system of two equations of degree four in two unknowns $\lambda_{1}$ and $\lambda_{2}$. This system has up to 16 real solutions and can be efficiently solved using the Gröbner basis method [23]. The final solver performs an elimination of a $20\times 36$ matrix and extracts solutions from the eigenvalues and eigenvectors of a $16\times 16$ matrix.
From up to 16 possible solutions, a solution $\lambda_{1}^{k}$ and $\lambda_{2}^{k}$ that minimizes $|\lambda_{1}|+|\lambda_{2}|$ is selected [26]. This solution is used to compute the $k^{th}$ iteration updates $\Delta f_{i}^{k}$ and $\Delta\mathbf{c}_{i}^{k}$ using equations (8)–(9) and subsequently the new estimates of focal lengths and principal points $s^{k}=\langle f_{1}^{k},f_{2}^{k},\mathbf{c}_{1}^{k},\mathbf{c}_{2}^{k}\rangle$.

### v3

@@ 1 Introduction
Camera calibration, i.e., the estimation of camera intrinsic parameters, is a fundamental problem in computer vision with many applications. The precision of estimated intrinsic parameters, such as focal length and principal point, significantly affects the precision of tasks in structure-from-motion (SfM) [snavely2008modeling],visual localization [Sattler-PAMI-2017], 3D object detection [xu_3dobj], augmented reality [Castle08ISWC], and other applications [sochor_vehiclespeed].
Classical calibration methods [zhang_calibration] use known calibration patterns, e.g. checkerboards, or additional knowledge of the observed scene to estimate the camera intrinsics. As such, they are often impractical. On the other hand, self-calibration methods do not require any additional knowledge of the scene geometry and rely on automatic detection of image features in input cameras. Thus, they are preferable in many applications.
@@
The geometry of two uncalibrated cameras is captured by the fundamental matrix [fundamental_matrix]. In 1998 Bougnoux [bougnoux_original] showed that assuming square pixels and known principal points of both cameras, it is possible to obtain the focal lengths of the two cameras from the fundamental matrix using a closed-form formula. This provides a convenient way to determine the focal length, since the principal point can usually be assumed to lie in the image center. However, the precision of the estimated focal lengths using the Bougnoux formula [bougnoux_original] and other similar methods [kanatani_closed_form, melanitis, fetzer] is often marred by inaccuracies, sometimes even resulting in physically implausible imaginary focal lengths. This stems from singularities and susceptibility of these methods to noise in fundamental matrices and assumed positions of principal points. Moreover, the singularities occur in very common camera configurations when the principal axes of the cameras are coplanar (i.e. they intersect), which is common when images of a single object of interest are taken from different views.
An alternative approach [hartley_iterative] relies on iterative optimization of a multiterm loss, allowing one to estimate both the focal lengths of the cameras as well as their principal points. This approach addresses some of the drawbacks of the aforementioned methods but still relies on the Bougnoux formula [bougnoux_original] within the iterative process. It is thus susceptible to inaccuracies, especially in the aforementioned degenerate camera configurations.
Similarly to [hartley_iterative], we formulate the problem of estimating focal lengths and principal points from a given fundamental matrix as a constrained optimization problem using priors. Constraints ensure that the solution satisfies the Kruppa equations [lourakis_calib]. To solve the constrained optimization problem, we search for all stationary points of its Lagrangian. This results in a complex system of polynomial equations that we efficiently solve in an iterative way. Each iteration in our formulation involves solving a system of two polynomial equations of degree four in two variables, which is efficiently solved using the Gröbner basis method [larsson_efficient].
Existing approaches, as well as our method, are based on fundamental matrices that can be estimated using different variants of RANSAC [ransac, magsac++, lo-ransac, degensac, vsac, prosac]. Some of the matrices produced by the 7-point algorithm [HZ-2003] within RANSAC lead to imaginary focal lengths when decomposed using the Bougnoux formula [bougnoux_original]. We observed that such matrices seldom lead to the final model. Based on this insight, we propose to perform a degeneracy check for such matrices and reject them, thus saving computational time on their further processing, such as scoring and local optimization.
@@
  • A novel iterative method for focal length and principal point estimation from fundamental matrices. Evaluation on synthetic and large-scale real-world datasets [phototourism_benchmark, aachen_dn_extended] shows that our method results in superior accuracy of estimated focal lengths as well as camera poses.
  • A simple and computationally efficient degeneracy check for fundamental matrices leading to imaginary focal lengths within RANSAC. Performance on real-world datasets [phototourism_benchmark, aachen_dn_extended] shows that performing this check and rejecting such models within various RANSAC variants [magsac++, lo-ransac, degensac, prosac] and implementations [poselib, opencv, vsac] leads to faster computation and more accurate pose and focal length estimates.
@@ 2.1 Direct Methods
Bougnoux [bougnoux_original] showed that under the assumption of square pixels and known positions of the principal points, it is possible to calculate the focal lengths of both cameras from the fundamental matrix in closed form. Kanatani and Matsunuga [kanatani_closed_form] derived a closed form solution directly from the elements of the fundamental matrix, avoiding the intermediate epipole computation required by the Bougnoux formula. Melanitis and Maragos [melanitis] formulated a linear system, the solution of which provides the focal lengths.
@@
Similar formulas have been derived for the case when the focal lengths of cameras are known to be equal [sturm_single, kanatani_closed_form, brooks_single]. In case of equal focal lengths, it is possible to avoid estimating the full 7-DoF fundamental matrix and instead employ a 6-point minimal solver [stewenius_single_6pt] within RANSAC to find the relative poses of cameras along with the focal length. These methods also suffer from generic singularities when the principal axes are parallel, or they intersect, and the camera centers are equidistant from the intersection point [sturm_degen].
An interesting case is when only one of the focal lengths is unknown. The focal length can be estimated in a closed form solution from the fundamental matrix [urbanek_onefocal] or using a 6-point minimal solver [bujnak_onefocal]. A degenerate case for these methods only occurs when the center of the calibrated camera lies on the principal axis of the uncalibrated camera, a situation rarely occurring in practice.
Since focal lengths appear in squared form in the Kruppa equations [lourakis_calib], all the previously mentioned direct methods provide only the squares of focal lengths. This may result in physically implausible imaginary focal lengths due to noise. Depending on the level of noise and camera configuration, imaginary focal lengths may occur at a significant rate (over 20% of samples) [kanatani_stabilizing]. Imaginary focal lengths are less likely to occur when self-calibration is performed using three views [kanatani_threeview].
@@ 2.2 Iterative Methods
Hartley and Silpa-Anan [hartley_iterative] pointed out that errors in the assumed positions of principal points may significantly affect the focal lengths estimated by the Bougnoux formula. To remedy this issue, they proposed an iterative method based on Levenberg-Marquardt optimization [lm]. During the optimization the fundamental matrix, along with the positions of principal points in both images, are optimized to minimize a multiterm loss. The loss consists of the Sampson error on the correspondences and the distance of the principal points and squared focal lengths from predetermined priors. To prevent imaginary focal lengths, the loss also contains a penalization term which dramatically increases the loss whenever the square focal lengths are below a given threshold. During the optimization procedure, the focal lengths are estimated from the fundamental matrix using the Bougnoux formula. This approach overcomes many of the issues of using the formula alone but may still suffer in the vicinity of degenerate camera configurations, as it still relies on the formula to estimate the focal lengths. Additionally, the procedure is computationally expensive as it requires computation of the Sampson error for all correspondences in each iteration.
Iterative methods can also be used to estimate the camera intrinsics using two or more views [gherardi_calib, pollefeys_calib, lourakis_calib, fetzer]. When considering only two views with two unknown focal lengths, these methods should converge to the Bougnoux formula. However, these methods may result in different outcomes in the vicinity of degenerate configurations. For example, the method by Fetzer et al[fetzer] optimizes for an energy functional based on the Kruppa equations [lourakis_calib]. In degenerate configurations, the functional does not have a single minimum, yet the optimization procedure will be terminated at some point, providing an output different from the one provided by the Bougnoux formula.
@@ 3 Robust focal length estimation
In this paper, similar to Hartley [hartley_iterative], we formulate the problem of estimating focal lengths from a given fundamental matrix $\mathbf{F}$ as an optimization problem using priors.
@@
For fixed principal points, e.g., $\mathbf{c}_{i}=[0,0]$, $i=1,2$, there is a closed form solution for the focal lengths extracted from the given fundamenatal matrix $\mathbf{F}$, i.e., the Bougnoux formula [bougnoux_original]. Since this formula can result in unstable or imaginary estimates, in our formulation, we allow moving the estimated principal points to avoid such instabilities. For non-fixed principal points $\mathbf{c}_{i}$, there are infinitely many decompositions of $\mathbf{F}$ into the essential matrix $\mathbf{E}$ and two calibration matrices of the form (1). Thus, we formulate the problem of estimating focal lengths from $\mathbf{F}$ as a constrained optimization, where during the optimization, the focal lengths, along with the positions of principal points in both images are optimized to minimize the following cost function:
@@
$\textrm{s.t.}\quad\kappa_{1}=$ $\quad\sigma_{1}(\mathbf{v}_{1}^{\top}\omega_{1}^{*}\mathbf{v}_{1})(\mathbf{u}_{1}^{\top}\omega_{2}^{*}\mathbf{u}_{2})+$
@@
where $f_{i}^{p}$ and $\mathbf{c}_{i}^{p}=[u_{i}^{p},v_{i}^{p}]^{\top}$, $i=1,2$ are the priors for the focal lengths and the principal points, $w_{i}^{f}$ and $w_{i}^{\mathbf{c}}$ are predetermined weights, and $\kappa_{1}=0$ and $\kappa_{2}=0$ are two Kruppa equations [lourakis_calib] [^2] Note, that there are three Kruppa equations, but only two are linearly independent., which are functions of both focal lengths and principal points that appear in $\omega^{*}_{i}=\mathbf{K}_{i}\mathbf{K}_{i}^{\top}$. The Kruppa equations ensure that for the input fundamental matrix $\mathbf{F}$ and the estimated calibration matrices $\mathbf{K}_{1}$ and $\mathbf{K}_{2}$ (1), the matrix $\mathbf{K}_{1}^{\top}\mathbf{F}\mathbf{K}_{2}$ is a valid essential matrix.
The cost function in (2) has several advantages over the cost function used in [hartley_iterative]. (1) It does not contain a term with the Sampson error on correspondences, which significantly slows down the optimization in [hartley_iterative]. (2) It operates directly on focal lengths $f_{i}$ instead of squared focal lengths $f_{i}^{2}$. (3) The minimization of (2) does not require computation or initialization using the Bougnoux formula [bougnoux_original] as used in [hartley_iterative]. Such an initialization can be very far from the ground truth focal lengths and in many scenarios can result in $f_{i}^{2}\leq 0$. Therefore, minimization of (2) does not need a special penalty term for $f_{i}^{2}\leq 0$ as used in [hartley_iterative].
@@
where $\lambda_{1}$ and $\lambda_{2}$ are the Lagrange multipliers. The constant ‘2-2’ is introduced for easier subsequent manipulation of the equations and it does not influence the final solution.
@@
Unfortunately, the system of eight equations (4)-(7) is too complex to be efficiently solved using an algebraic method, e.g. using the Gröbner basis method [kukelova_automatic, larsson_efficient]. Next, we propose an iterative method to efficiently solve this system of polynomial equations. Note that in our method, we estimate all stationary points of the Lagrange function $L$ (3). Thus, we avoid potential problems of numerical methods that directly minimize (3). Such methods usually require good initialization and can get stuck in local minima.
@@ 3.1 Iterative method
The proposed iterative solution to equations (4)–(7) follows the idea of iterative triangulation methods [Lindstrom-CVPR-2010, Kukelova_rad_triang]. First, let $s^{k-1}=\langle f_{1}^{k-1},f_{2}^{k-1},\mathbf{c}_{1}^{k-1},\mathbf{c}_{2}^{k-1}\rangle$ denote the best current estimate of focal lengths and principal points after the $(k-1)^{th}$ iteration. The prior values $f_{i}^{0}\equiv f_{i}^{p}$, $\mathbf{c}_{i}^{0}\equiv\mathbf{c}_{i}^{p}$, $i=1,2$ are used as an initialization. In the $k^{th}$ iteration the updated estimates $s^{k}=\langle f_{1}^{k},f_{2}^{k},\mathbf{c}_{1}^{k},\mathbf{c}_{2}^{k}\rangle$ are obtained by replacing the focal lengths and principal points $\langle f_{1},f_{2},\mathbf{c}_{1},\mathbf{c}_{2}\rangle$ on the right-hand side of equations (8)–(9) by the current best estimates $s^{k-1}=\langle f_{1}^{k-1},f_{2}^{k-1},\mathbf{c}_{1}^{k-1},\mathbf{c}_{2}^{k-1}\rangle$. In this way, we obtain the expressions for $\Delta f_{i}=\frac{1}{w_{i}^{f}}(\lambda_{1}\frac{\partial\kappa_{1}}{\partial f_{i}}(s^{k-1})+\lambda_{2}\frac{\partial\kappa_{2}}{\partial f_{i}}(s^{k-1}))$ and $\Delta\mathbf{c}_{i}=\frac{1}{w_{i}^{\mathbf{c}}}(\lambda_{1}(\frac{\partial\kappa_{1}}{\partial\mathbf{c}_{i}}(s^{k-1}))^{\top}+\lambda_{2}(\frac{\partial\kappa_{2}}{\partial\mathbf{c}_{i}}(s^{k-1}))^{\top})$, $i=1,2$. Here $\frac{\partial\kappa_{j}}{\partial f_{i}}(s^{k-1})$ and $\frac{\partial\kappa_{j}}{\partial\mathbf{c}_{i}}(s^{k-1})$, $i,j=1,2$ are partial derivatives of $\kappa_{j}$ evaluated at $s^{k-1}=\langle f_{1}^{k-1},f_{2}^{k-1},\mathbf{c}_{1}^{k-1},\mathbf{c}_{2}^{k-1}\rangle$. After this substitution, $\Delta f_{i}$ and $\Delta\mathbf{c}_{i}$ are functions of two Lagrange multipliers $\lambda_{1}$ and $\lambda_{2}$.
@@
The updated Kruppa equations (6) and (7), which we denote $\kappa_{1}^{k}=0$ and $\kappa_{2}^{k}=0$, form a system of two equations of degree four in two unknowns $\lambda_{1}$ and $\lambda_{2}$. This system has up to 16 real solutions and can be efficiently solved using the Gröbner basis method [larsson_efficient]. The final solver performs an elimination of a $20\times 36$ matrix and extracts solutions from the eigenvalues and eigenvectors of a $16\times 16$ matrix.
From up to 16 possible solutions, a solution $\lambda_{1}^{k}$ and $\lambda_{2}^{k}$ that minimizes $|\lambda_{1}|+|\lambda_{2}|$ is selected [Lindstrom-CVPR-2010]. This solution is used to compute the $k^{th}$ iteration updates $\Delta f_{i}^{k}$ and $\Delta\mathbf{c}_{i}^{k}$ using equations (8)–(9) and subsequently the new estimates of focal lengths and principal points $s^{k}=\langle f_{1}^{k},f_{2}^{k},\mathbf{c}_{1}^{k},\mathbf{c}_{2}^{k}\rangle$.